OpenStudy (anonymous):

Compute d2y/dx2 at the point (4, 3). x^2-y^2 = 7

OpenStudy (anonymous):

dx^2/dx - dy^2/dx = d7/dx = 0. So, 2x-(dy^2/dy)(dy/dx)=0, 2x-2y(dy/dx)=0, x-y(dy/dx)=0, x=y(dy/dx), x/y=dy/dx, dy/dx=x/y, d^2y/dx^2=d(x/y)/dx=[(dx/dx)y-x(dy/dx)]/y^2, d^2y/dx^2=[1y-x(dy/dx)]/y^2, substituing dy/dx by x/y d^2y/dx^2=[1y-x^2y]/y^2, in the point(4,3), x=4 and y=3. So, d^2y/dx^2=[3-(4^2)3]/3^2, Then, d^2y/dx^2=[3-48]/9=-5

OpenStudy (anonymous):

there are answer choices

OpenStudy (anonymous):

which one is it? ill post it now

OpenStudy (anonymous):

a) 7/27 b)- 7/27 c) 14/ 27 d) - 14/27 e) - 7/9

OpenStudy (anonymous):

oh sorry, in the 3rd line from down to up the coorect is d^2y/dx^2=[1y-(x^2)/y]/y^2, so substituing d^2y/dx^2=[3-(4^2)/3]/(3^2). Then, d^2y/dx^2=-7/27, letterb

OpenStudy (anonymous):

thank you....i was kinda confused y u got five wen its not in any of the answer choices

OpenStudy (anonymous):

no problem